# Environmentally Sustainable Software Design and Development: A Systematic Literature Review

### v1

@@ Acknowledgements.
- Goaër (2021) Olivier Le Goaër. 2021. Enforcing green code with Android lint. In Proceedings of the 35th IEEE/ACM International Conference on Automated Software Engineering. Association for Computing Machinery, New York, NY, USA, 85–90. https://doi.org/10.1145/3417113.3422188
- Goaër (2023) Olivier Le Goaër. 2023. Decarbonizing Software with Free and Open Source Software: The ecoCode Project. In Proceedings of the Research Projects Exhibition Papers Presented at the 35th International Conference on Advanced Information Systems Engineering (CAiSE 2023), Zaragoza, Spain, June 12-16, 2023 (CEUR Workshop Proceedings, Vol. 3413), Jaime Font, Lorena Arcega, José Fabián Reyes Román, and Giovanni Giachetti (Eds.). CEUR-WS.org, 9–13. https://ceur-ws.org/Vol-3413/paper2.pdf
- Goaër and Hertout (2022) Olivier Le Goaër and Julien Hertout. 2022. ecoCode: a SonarQube Plugin to Remove Energy Smells from Android Projects. In 37th IEEE/ACM International Conference on Automated Software Engineering, ASE 2022, Rochester, MI, USA, October 10-14, 2022. ACM, 157:1–157:4. https://doi.org/10.1145/3551349.3559518
- Heithoff et al. (2023) Malte Heithoff, Alexander David Hellwig, Judith Michael, and Bernhard Rumpe. 2023. Digital Twins for Sustainable Software Systems. In 7th IEEE/ACM International Workshop on Green And Sustainable Software, GREENS@ICSE 2023, Melbourne, Australia, May 14, 2023. IEEE, 19–23. https://doi.org/10.1109/GREENS59328.2023.00010
- Hoesch-Klohe et al. (2010) Konstantin Hoesch-Klohe, Aditya Ghose, and Lam-Son Lê. 2010. Towards Green Business Process Management. In 2010 IEEE International Conference on Services Computing, SCC 2010, Miami, Florida, USA, July 5-10, 2010. IEEE Computer Society, 386–393. https://doi.org/10.1109/SCC.2010.21
- Ibrahim et al. (2021) Siti Rohana Ahmad Ibrahim, Jamaiah Yahaya, Hasimi Salehudin, and Aziz Deraman. 2021. The Development of Green Software Process Model: A Qualitative Design and Pilot Study. International Journal of Advanced Computer Science and Applications 12, 8 (2021), 589 – 598. https://doi.org/10.14569/IJACSA.2021.0120869
@@
- Jayanthi et al. (2021) Harika Jayanthi, Anjali Kulkarni, Aishwarya E Patil, and Sheela S V. 2021. An Organizational Structure for Sustainable Software Development. In Proceedings - 2021 3rd International Conference on Advances in Computing, Communication Control and Networking, ICAC3N 2021. 1539 – 1542. https://doi.org/10.1109/ICAC3N53548.2021.9725722
@@
- Johann et al. (2011) Timo Johann, Markus Dick, Eva Kern, and Stefan Naumann. 2011. Sustainable development, sustainable software, and sustainable software engineering: An integrated approach. In 2011 International Symposium on Humanities, Science and Engineering Research. 34–39. https://doi.org/10.1109/SHUSER.2011.6008495
- Karunakaran and Rao (2013) Deepak Karunakaran and G Subrahmanya Vrk Rao. 2013. A petri net simulation of software development lifecycle towards green IT. In 2013 IEEE Conference on Open Systems, ICOS 2013. 58 – 62. https://doi.org/10.1109/ICOS.2013.6735048
@@
- Kipp et al. (2011) Alexander Kipp, Tao Jiang, and Mariagrazia Fugini. 2011. Green Metrics for Energy-aware IT Systems. In International Conference on Complex, Intelligent and Software Intensive Systems, CISIS 2011, June 30 - July 2, 2011, Korean Bible University, Seoul, Korea. IEEE Computer Society, 241–248. https://doi.org/10.1109/CISIS.2011.42
@@
- Koçak and Alptekin (2019) Sedef Akinli Koçak and Gülfem Isiklar Alptekin. 2019. A Utility Model for Designing Environmentally Sustainable Software. In Proceedings of the 8th International Workshop on Requirements Engineering for Sustainable Systems co-located with 27th International Conference on Requirements Engineering, Jeju, South Korea, September 24, 2019 (CEUR Workshop Proceedings, Vol. 2541), Colin C. Venters Birgit Penzenstadler and Ruzanna Chitchyan (Eds.). CEUR-WS.org. https://ceur-ws.org/Vol-2541/paper8.pdf
- Koziolek (2011) Heiko Koziolek. 2011. Sustainability evaluation of software architectures: a systematic review. In 7th International Conference on the Quality of Software Architectures, QoSA 2011 and 2nd International Symposium on Architecting Critical Systems, ISARCS 2011. Boulder, CO, USA, June 20-24, 2011, Proceedings, Ivica Crnkovic, Judith A. Stafford, Dorina C. Petriu, Jens Happe, and Paola Inverardi (Eds.). ACM, 3–12. https://doi.org/10.1145/2000259.2000263
@@
- Lago (2019) Patricia Lago. 2019. Architecture design decision maps for software sustainability. In Proceedings of the 41st International Conference on Software Engineering: Software Engineering in Society, ICSE 2019, Montreal, QC, Canada, May 25-31, 2019, Rick Kazman and Liliana Pasquale (Eds.). ACM, 61–64. https://doi.org/10.1109/ICSE-SEIS.2019.00015
- Lago and Jansen (2010) Patricia Lago and Toon Jansen. 2010. Creating Environmental Awareness in Service Oriented Software Engineering. In Service-Oriented Computing - ICSOC 2010 International Workshops, PAASC, WESOA, SEE, and SOC-LOG, San Francisco, CA, USA, December 7-10, 2010, Revised Selected Papers (Lecture Notes in Computer Science, Vol. 6568), E. Michael Maximilien, Gustavo Rossi, Soe-Tsyr Yuan, Heiko Ludwig, and Marcelo Fantinato (Eds.). 181–186. https://doi.org/10.1007/978-3-642-19394-1_19
- Lami and Buglione (2012) Giuseppe Lami and Luigi Buglione. 2012. Measuring Software Sustainability from a Process-centric Perspective. In 2012 Joint Conference of the 22nd International Workshop on Software Measurement and the 2012 Seventh International Conference on Software Process and Product Measurement, Assisi, Italy, October 17-19, 2012. IEEE Computer Society, 53–59. https://doi.org/10.1109/IWSM-MENSURA.2012.16
- Lami et al. (2014) Giuseppe Lami, Fabrizio Fabbrini, and Luigi Buglione. 2014. An ISO/IEC 33000-Compliant Measurement Framework for Software Process Sustainability Assessment. In 2014 Joint Conference of the International Workshop on Software Measurement and the International Conference on Software Process and Product Measurement, Rotterdam, The Netherlands, October 6-8, 2014. IEEE Computer Society, 50–59. https://doi.org/10.1109/IWSM.MENSURA.2014.34
- Lami et al. (2013) Giuseppe Lami, Fabrizio Fabbrini, and Mario Fusani. 2013. A methodology to derive sustainability indicators for software development projects. In International Conference on Software and System Process, ICSSP ’13, San Francisco, CA, USA, May 18-19, 2013, Jürgen Münch, Jo Ann Lan, and He Zhang (Eds.). ACM, 70–77. https://doi.org/10.1145/2486046.2486060
@@
- Mahmoud and Ahmad (2013) Sara Mahmoud and Imtiaz Ahmad. 2013. A Green Model for Sustainable Software Engineering. International Journal of Software Engineering and Its Applications 7, 4 (2013).
- Mancebo et al. (2021) Javier Mancebo, Coral Calero, Félix García, María Ángeles Moraga, and Ignacio García Rodríguez de Guzmán. 2021. FEETINGS: Framework for Energy Efficiency Testing to Improve eNvironmental Goals of the Software. In Software Sustainability, Coral Calero, María Ángeles Moraga, and Mario Piattini (Eds.). Springer, 69–87. https://doi.org/10.1007/978-3-030-69970-3_4
- Mehra et al. (2023) Rohit Mehra, Priyavanshi Pathania, Vibhu Saujanya Sharma, Vikrant Kaulgud, Sanjay Podder, and Adam P. Burden. 2023. Assessing the Impact of Refactoring Energy-Inefficient Code Patterns on Software Sustainability: An Industry Case Study. In 38th IEEE/ACM International Conference on Automated Software Engineering, ASE 2023, Luxembourg, September 11-15, 2023. IEEE, 1825–1827. https://doi.org/10.1109/ASE56229.2023.00205
- Mehra et al. (2022) Rohit Mehra, Vibhu Saujanya Sharma, Vikrant Kaulgud, Sanjay Podder, and Adam P. Burden. 2022. Towards a Green Quotient for Software Projects. In 44th IEEE/ACM International Conference on Software Engineering: Software Engineering in Practice, ICSE (SEIP) 2022, Pittsburgh, PA, USA, May 22-24, 2022. IEEE, 295–296. https://doi.org/10.1109/ICSE-SEIP55303.2022.9793921
@@
- Moshnyaga (2013a) Vasily G. Moshnyaga. 2013a. An assessment of software lifecycle energy. In 2013 23rd International Workshop on Power and Timing Modeling, Optimization and Simulation (PATMOS), Karlsruhe, Germany, September 9-11, 2013. IEEE, 112–119. https://doi.org/10.1109/PATMOS.2013.6662163
- Moshnyaga (2013b) Vasily G. Moshnyaga. 2013b. Assessment of Software Lifecycle Energy and its Contribution to Green House Gas Emissions. In 2013 International Conference on Cloud and Green Computing, Karlsruhe, Germany, September 30 - October 2, 2013. IEEE Computer Society, 197–198. https://doi.org/10.1109/CGC.2013.37
- Moshnyaga (2013c) Vasily G. Moshnyaga. 2013c. Lifecycle energy assessment of mobile applications. In Proceedings of the 2013 International Workshop on Software Development Lifecycle for Mobile, DeMobile 2013, Saint Petersburg, Russia, August 19, 2013, Aharon Abadi, Rafael Prikladnicki, and Yael Dubinsky (Eds.). ACM, 17–23. https://doi.org/10.1145/2501553.2501557
@@
- Noureddine et al. (2012) Adel Noureddine, Aurelien Bourdon, Romain Rouvoy, and Lionel Seinturier. 2012. A preliminary study of the impact of software engineering on GreenIT. In First International Workshop on Green and Sustainable Software, GREENS 2012, Zurich, Switzerland, June 3, 2012, Rick Kazman, Patricia Lago, Niklaus Meyer, Maurizio Morisio, Hausi A. Müller, Frances Paulisch, Giuseppe Scanniello, and Olaf Zimmermann (Eds.). IEEE Computer Society, 21–27. https://doi.org/10.1109/GREENS.2012.6224251
@@
- Ormazabal et al. (2014) Marta Ormazabal, Carmen Jaca, and Rogério Puga-Leal. 2014. Analysis and Comparison of Life Cycle Assessment and Carbon Footprint Software. In Proceedings of the Eighth International Conference on Management Science and Engineering Management, Jiuping Xu, Virgílio António Cruz-Machado, Benjamin Lev, and Stefan Nickel (Eds.). Springer Berlin Heidelberg, Berlin, Heidelberg, 1521–1530.
- Pa et al. (2017) Noraini Che Pa, Faizal Karim, and Sa’adah Hassan. 2017. Dashboard system for measuring green software design. In Proceeding - 2017 3rd International Conference on Science in Information Technology: Theory and Application of IT for Education, Industry and Society in Big Data Era, ICSITech 2017, Vol. 2018-January. 325 – 329. https://doi.org/10.1109/ICSITech.2017.8257133
- Pan and Rajan (2022) Rangeet Pan and Hridesh Rajan. 2022. Decomposing convolutional neural networks into reusable and replaceable modules. In Proceedings of the 44th International Conference on Software Engineering (Pittsburgh, Pennsylvania) (ICSE ’22). Association for Computing Machinery, New York, NY, USA, 524–535. https://doi.org/10.1145/3510003.3510051
@@
- Penzenstadler et al. (2012) Birgit Penzenstadler, Veronika Bauer, Coral Calero, and Xavier Franch. 2012. Sustainability in software engineering: A systematic literature review. In 16th International Conference on Evaluation & Assessment in Software Engineering, EASE 2012, Ciudad Real, Spain, May 14-15, 2012. Proceedings, Maria Teresa Baldassarre, Marcela Genero, Emilia Mendes, and Mario Piattini (Eds.). IET - The Institute of Engineering and Technology / IEEE Xplore, 32–41. https://doi.org/10.1049/IC.2012.0004
- Penzenstadler and Femmer (2013) Birgit Penzenstadler and Henning Femmer. 2013. A generic model for sustainability with process- and product-specific instances. In Proceedings of the 2013 Workshop on Green in/by Software Engineering, GIBSE ’13, Fukuoka, Japan, March 26, 2013, Somayeh Malakuti, Christoph Bockisch, Uwe Assmann, Sebastian Götz, Steven te Brinke, Birgit Penzenstadler, and Birgit Demuth (Eds.). ACM, 3–8. https://doi.org/10.1145/2451605.2451609
@@
- Ponsard et al. (2018) Christophe Ponsard, Renaud De Landtsheer, and Fabian Germeau. 2018. Building sustainable software for sustainable systems: case study of a shared pick-up and delivery service. In Proceedings of the 6th International Workshop on Green and Sustainable Software, GREENS@ICSE 2018, Gothenburg, Sweden, May 27, 2018, Ivano Malavolta, Rick Kazman, and João Saraiva (Eds.). ACM, 31–34. https://doi.org/10.1145/3194078.3194083
@@
- Raisian et al. (2021) Komeil Raisian, Jamaiah Yahaya, Siti Rohana Ahmad Ibrahim, Aziz Deraman, and Tumen Yunos. 2021. The Green Software Measurement Structure Based on Sustainability Perspective. In International Conference on Electrical Engineering and Informatics, ICEEI 2021, Kuala Terengganu, Malaysia, October 12-13, 2021. IEEE, 1–6. https://doi.org/10.1109/ICEEI52609.2021.9611108
@@
- Sharma et al. (2015) Sunil Kumar Sharma, P. K. Gupta, and Reza Malekian. 2015. Energy efficient software development life cycle - An approach towards smart computing. In 2015 IEEE International Conference on Computer Graphics, Vision and Information Security (CGVIS). 1–5. https://doi.org/10.1109/CGVIS.2015.7449881
- Shenoy and Eeratta (2011) Sanath. S. Shenoy and Raghavendra Eeratta. 2011. Green software development model: An approach towards sustainable software development. In 2011 Annual IEEE India Conference. 1–6. https://doi.org/10.1109/INDCON.2011.6139638
@@
- Sikand et al. (2023) Samarth Sikand, Vibhu Saujanya Sharma, Vikrant Kaulgud, and Sanjay Podder. 2023. Green AI Quotient: Assessing Greenness of AI-based software and the way forward. In 38th IEEE/ACM International Conference on Automated Software Engineering, ASE 2023, Luxembourg, September 11-15, 2023. IEEE, 1828–1833. https://doi.org/10.1109/ASE56229.2023.00115
- Simon et al. (2023) Thibault Simon, Pierre Rust, Romain Rouvoy, and Joël Penhoat. 2023. Uncovering the Environmental Impact of Software Life Cycle. In International Conference on ICT for Sustainability, ICT4S 2023, Rennes, France, June 5-9, 2023. IEEE, 176–187. https://doi.org/10.1109/ICT4S58814.2023.00026
@@
- Tee et al. (2014a) Mcxin Tee, Rusli Abdullah, and Salfarina Abdullah. 2014a. Toward Developing Green Software Development Model in Managing Knowledge of IT Practitioners for Sustaining Future Generation. In Knowledge Management International Conference 2014 (KMICe).
- Tee et al. (2014b) Mcxin Tee, Rusli Abdullah, and Salfarina Abdullah. 2014b. Web-based knowledge management model for managing and sharing green knowledge of software development in community of practice. In 2014 8th. Malaysian Software Engineering Conference (MySEC). 210–215. https://doi.org/10.1109/MySec.2014.6986016
@@
- Verdecchia et al. (2021) Roberto Verdecchia et al. 2021. Green IT and Green Software. IEEE Softw. 38, 6 (2021), 7–15. https://doi.org/10.1109/MS.2021.3102254

### v2

@@ Acknowledgements.
- Goaër (2021) Olivier Le Goaër. 2021. Enforcing green code with Android lint. In Proceedings of the 35th IEEE/ACM Int. Conf. on Automated Software Engineering. Association for Computing Machinery, New York, NY, USA, 85–90. https://doi.org/10.1145/3417113.3422188
- Goaër (2023) Olivier Le Goaër. 2023. Decarbonizing Software with Free and Open Source Software: The ecoCode Project. In Proceedings of the Research Projects Exhibition Papers Presented at the 35th Int. Conf. on Advanced Information Systems Engineering (CAiSE 2023), Zaragoza, Spain, June 12-16, 2023 (CEUR Workshop Proceedings, Vol. 3413), Jaime Font et al. (Eds.). CEUR-WS.org, 9–13. https://ceur-ws.org/Vol-3413/paper2.pdf
- Goaër and Hertout (2022) Olivier Le Goaër and Julien Hertout. 2022. ecoCode: a SonarQube Plugin to Remove Energy Smells from Android Projects. In 37th IEEE/ACM Int. Conf. on Automated Software Engineering, ASE 2022, Rochester, MI, USA, October 10-14, 2022. ACM, 157:1–157:4. https://doi.org/10.1145/3551349.3559518
- Heithoff et al. (2023) Malte Heithoff, Alexander David Hellwig, Judith Michael, and Bernhard Rumpe. 2023. Digital Twins for Sustainable Software Systems. In 7th IEEE/ACM Int. Workshop on Green And Sustainable Software, GREENS@ICSE 2023, Melbourne, Australia, May 14, 2023. IEEE, 19–23. https://doi.org/10.1109/GREENS59328.2023.00010
- Hoesch-Klohe et al. (2010) Konstantin Hoesch-Klohe, Aditya Ghose, and Lam-Son Lê. 2010. Towards Green Business Process Management. In 2010 IEEE Int. Conf. on Services Computing, SCC 2010, Miami, Florida, USA, July 5-10, 2010. IEEE Computer Society, 386–393. https://doi.org/10.1109/SCC.2010.21
- Ibrahim et al. (2021) Siti Rohana Ahmad Ibrahim, Jamaiah Yahaya, Hasimi Salehudin, and Aziz Deraman. 2021. The Development of Green Software Process Model: A Qualitative Design and Pilot Study. Int. Journal of Advanced Computer Science and Applications 12, 8 (2021), 589 – 598. https://doi.org/10.14569/IJACSA.2021.0120869
@@
- Jayanthi et al. (2021) Harika Jayanthi, Anjali Kulkarni, Aishwarya E Patil, and Sheela S V. 2021. An Organizational Structure for Sustainable Software Development. In Proceedings - 2021 3rd Int. Conf. on Advances in Computing, Communication Control and Networking, ICAC3N 2021. 1539 – 1542. https://doi.org/10.1109/ICAC3N53548.2021.9725722
@@
- Johann et al. (2011) Timo Johann, Markus Dick, Eva Kern, and Stefan Naumann. 2011. Sustainable development, sustainable software, and sustainable software engineering: An integrated approach. In 2011 Int. Symposium on Humanities, Science and Engineering Research. 34–39. https://doi.org/10.1109/SHUSER.2011.6008495
- Karunakaran and Rao (2013) Deepak Karunakaran and G Subrahmanya Vrk Rao. 2013. A petri net simulation of software development lifecycle towards green IT. In 2013 IEEE Conf. on Open Systems, ICOS 2013. 58 – 62. https://doi.org/10.1109/ICOS.2013.6735048
@@
- Kipp et al. (2011) Alexander Kipp, Tao Jiang, and Mariagrazia Fugini. 2011. Green Metrics for Energy-aware IT Systems. In Int. Conf. on Complex, Intelligent and Software Intensive Systems, CISIS 2011, June 30 - July 2, 2011, Korean Bible University, Seoul, Korea. IEEE Computer Society, 241–248. https://doi.org/10.1109/CISIS.2011.42
@@
- Koçak and Alptekin (2019) Sedef Akinli Koçak and Gülfem Isiklar Alptekin. 2019. A Utility Model for Designing Environmentally Sustainable Software. In Proceedings of the 8th Int. Workshop on Requirements Engineering for Sustainable Systems co-located with 27th Int. Conf. on Requirements Engineering, Jeju, South Korea, September 24, 2019 (CEUR Workshop Proceedings, Vol. 2541), Birgit Penzenstadler et al. (Eds.). CEUR-WS.org. https://ceur-ws.org/Vol-2541/paper8.pdf
- Koziolek (2011) Heiko Koziolek. 2011. Sustainability evaluation of software architectures: a systematic review. In 7th International Conference on the Quality of Software Architectures, QoSA 2011 and 2nd International Symposium on Architecting Critical Systems, ISARCS 2011. Boulder, CO, USA, June 20-24, 2011, Proceedings, Ivica Crnkovic et al. (Eds.). ACM, 3–12. https://doi.org/10.1145/2000259.2000263
@@
- Lago (2019) Patricia Lago. 2019. Architecture design decision maps for software sustainability. In Proceedings of the 41st Int. Conf. on Software Engineering: Software Engineering in Society, ICSE 2019, Montreal, QC, Canada, May 25-31, 2019, Rick Kazman and Liliana Pasquale (Eds.). ACM, 61–64. https://doi.org/10.1109/ICSE-SEIS.2019.00015
- Lago and Jansen (2010) Patricia Lago and Toon Jansen. 2010. Creating Environmental Awareness in Service Oriented Software Engineering. In Service-Oriented Computing - ICSOC 2010 Int. Workshops, PAASC, WESOA, SEE, and SOC-LOG, San Francisco, CA, USA, December 7-10, 2010, Revised Selected Papers (Lecture Notes in Computer Science, Vol. 6568), E. Michael Maximilien et al. (Eds.). 181–186. https://doi.org/10.1007/978-3-642-19394-1_19
- Lami and Buglione (2012) Giuseppe Lami and Luigi Buglione. 2012. Measuring Software Sustainability from a Process-centric Perspective. In 2012 Joint Conf. of the 22nd Int. Workshop on Software Measurement and the 2012 Seventh Int. Conf. on Software Process and Product Measurement, Assisi, Italy, October 17-19, 2012. IEEE Computer Society, 53–59. https://doi.org/10.1109/IWSM-MENSURA.2012.16
- Lami et al. (2014) Giuseppe Lami, Fabrizio Fabbrini, and Luigi Buglione. 2014. An ISO/IEC 33000-Compliant Measurement Framework for Software Process Sustainability Assessment. In 2014 Joint Conf. of the Int. Workshop on Software Measurement and the Int. Conf. on Software Process and Product Measurement, Rotterdam, The Netherlands, October 6-8, 2014. IEEE Computer Society, 50–59. https://doi.org/10.1109/IWSM.MENSURA.2014.34
- Lami et al. (2013) Giuseppe Lami, Fabrizio Fabbrini, and Mario Fusani. 2013. A methodology to derive sustainability indicators for software development projects. In Int. Conf. on Software and System Process, ICSSP ’13, San Francisco, CA, USA, May 18-19, 2013, Jürgen Münch et al. (Eds.). ACM, 70–77. https://doi.org/10.1145/2486046.2486060
@@
- Mahmoud and Ahmad (2013) Sara Mahmoud and Imtiaz Ahmad. 2013. A Green Model for Sustainable Software Engineering. Int. Journal of Software Engineering and Its Applications 7, 4 (2013).
- Mancebo et al. (2021) Javier Mancebo, Coral Calero, Félix García, María Ángeles Moraga, and Ignacio García Rodríguez de Guzmán. 2021. FEETINGS: Framework for Energy Efficiency Testing to Improve eNvironmental Goals of the Software. In Software Sustainability, Coral Calero et al. (Eds.). Springer, 69–87. https://doi.org/10.1007/978-3-030-69970-3_4
- Manner (2022) Jukka Manner. 2022. Black software — the energy unsustainability of software systems in the 21st century. Oxford Open Energy 2 (12 2022), oiac011. https://doi.org/10.1093/ooenergy/oiac011
- Mehra et al. (2023) Rohit Mehra, Priyavanshi Pathania, Vibhu Saujanya Sharma, Vikrant Kaulgud, Sanjay Podder, and Adam P. Burden. 2023. Assessing the Impact of Refactoring Energy-Inefficient Code Patterns on Software Sustainability: An Industry Case Study. In 38th IEEE/ACM Int. Conf. on Automated Software Engineering, ASE 2023, Luxembourg, September 11-15, 2023. IEEE, 1825–1827. https://doi.org/10.1109/ASE56229.2023.00205
- Mehra et al. (2022) Rohit Mehra, Vibhu Saujanya Sharma, Vikrant Kaulgud, Sanjay Podder, and Adam P. Burden. 2022. Towards a Green Quotient for Software Projects. In 44th IEEE/ACM Int. Conf. on Software Engineering: Software Engineering in Practice, ICSE (SEIP) 2022, Pittsburgh, PA, USA, May 22-24, 2022. IEEE, 295–296. https://doi.org/10.1109/ICSE-SEIP55303.2022.9793921
- Mingay and Pamlin (2008) Simon Mingay and Dennis Pamlin. 2008. Assessment of Global Low-Carbon and Environmental Leadership in the ICT Sector. Vol. ID G00161337, Gartner and WWF.
@@
- Möller et al. (2024) Tessa Möller, Annika Ernest Högner, Carl-Friedrich Schleussner, Samuel Bien, Niklas H. Kitzmann, Robin D. Lamboll, Joeri Rogelj, Jonathan F. Donges, Johan Rockström, and Nico Wunderling. 2024. Achieving net zero greenhouse gas emissions critical to limit climate tipping risks. Nature Communications 15 (2024). Issue 1. https://doi.org/10.1038/s41467-024-49863-0
@@
- Moshnyaga (2013a) Vasily G. Moshnyaga. 2013a. An assessment of software lifecycle energy. In 2013 23rd Int. Workshop on Power and Timing Modeling, Optimization and Simulation (PATMOS), Karlsruhe, Germany, September 9-11, 2013. IEEE, 112–119. https://doi.org/10.1109/PATMOS.2013.6662163
- Moshnyaga (2013b) Vasily G. Moshnyaga. 2013b. Assessment of Software Lifecycle Energy and its Contribution to Green House Gas Emissions. In 2013 Int. Conf. on Cloud and Green Computing, Karlsruhe, Germany, September 30 - October 2, 2013. IEEE Computer Society, 197–198. https://doi.org/10.1109/CGC.2013.37
- Moshnyaga (2013c) Vasily G. Moshnyaga. 2013c. Lifecycle energy assessment of mobile applications. In Proceedings of the 2013 Int. Workshop on Software Development Lifecycle for Mobile, DeMobile 2013, Saint Petersburg, Russia, August 19, 2013, Aharon Abadi et al. (Eds.). ACM, 17–23. https://doi.org/10.1145/2501553.2501557
@@
- Murugesan (2021) San Murugesan. 2021. 10 Recommendations for Green Software Development. Green Software Foundation, https://greensoftware.foundation/articles/10-recommendations-for-green-software-development.
@@
- Noureddine et al. (2012) Adel Noureddine, Aurelien Bourdon, Romain Rouvoy, and Lionel Seinturier. 2012. A preliminary study of the impact of software engineering on GreenIT. In First Int. Workshop on Green and Sustainable Software, GREENS 2012, Zurich, Switzerland, June 3, 2012, Rick Kazman et al. (Eds.). IEEE Computer Society, 21–27. https://doi.org/10.1109/GREENS.2012.6224251
@@
- Ormazabal et al. (2014) Marta Ormazabal, Carmen Jaca, and Rogério Puga-Leal. 2014. Analysis and Comparison of Life Cycle Assessment and Carbon Footprint Software. In Proceedings of the Eighth International Conference on Management Science and Engineering Management, Jiuping Xu et al. (Eds.). Springer Berlin Heidelberg, Berlin, Heidelberg, 1521–1530.
- Pa et al. (2017) Noraini Che Pa, Faizal Karim, and Sa’adah Hassan. 2017. Dashboard system for measuring green software design. In Proceeding - 2017 3rd Int. Conf. on Science in Information Technology: Theory and Application of IT for Education, Industry and Society in Big Data Era, ICSITech 2017, Vol. 2018-January. 325 – 329. https://doi.org/10.1109/ICSITech.2017.8257133
- Pan and Rajan (2022) Rangeet Pan and Hridesh Rajan. 2022. Decomposing convolutional neural networks into reusable and replaceable modules. In Proceedings of the 44th Int. Conf. on Software Engineering (Pittsburgh, Pennsylvania) (ICSE ’22). Association for Computing Machinery, New York, NY, USA, 524–535. https://doi.org/10.1145/3510003.3510051
@@
- Penzenstadler et al. (2012) Birgit Penzenstadler, Veronika Bauer, Coral Calero, and Xavier Franch. 2012. Sustainability in software engineering: A systematic literature review. In 16th International Conference on Evaluation & Assessment in Software Engineering, EASE 2012, Ciudad Real, Spain, May 14-15, 2012. Proceedings, Maria Teresa Baldassarre et al. (Eds.). IET / IEEE Xplore, 32–41. https://doi.org/10.1049/IC.2012.0004
- Penzenstadler and Femmer (2013) Birgit Penzenstadler and Henning Femmer. 2013. A generic model for sustainability with process- and product-specific instances. In Proceedings of the 2013 Workshop on Green in/by Software Engineering, GIBSE ’13, Fukuoka, Japan, March 26, 2013, Somayeh Malakuti et al. (Eds.). ACM, 3–8. https://doi.org/10.1145/2451605.2451609
@@
- Ponsard et al. (2018) Christophe Ponsard, Renaud De Landtsheer, and Fabian Germeau. 2018. Building sustainable software for sustainable systems: case study of a shared pick-up and delivery service. In Proceedings of the 6th Int. Workshop on Green and Sustainable Software, GREENS@ICSE 2018, Gothenburg, Sweden, May 27, 2018, Ivano Malavolta et al. (Eds.). ACM, 31–34. https://doi.org/10.1145/3194078.3194083
@@
- Raisian et al. (2021) Komeil Raisian, Jamaiah Yahaya, Siti Rohana Ahmad Ibrahim, Aziz Deraman, and Tumen Yunos. 2021. The Green Software Measurement Structure Based on Sustainability Perspective. In Int. Conf. on Electrical Engineering and Informatics, ICEEI 2021, Kuala Terengganu, Malaysia, October 12-13, 2021. IEEE, 1–6. https://doi.org/10.1109/ICEEI52609.2021.9611108
@@
- Sharma et al. (2015) Sunil Kumar Sharma, P. K. Gupta, and Reza Malekian. 2015. Energy efficient software development life cycle - An approach towards smart computing. In 2015 IEEE Int. Conf. on Computer Graphics, Vision and Information Security (CGVIS). 1–5. https://doi.org/10.1109/CGVIS.2015.7449881
- Shenoy and Eeratta (2011) Sanath. S. Shenoy and Raghavendra Eeratta. 2011. Green software development model: An approach towards sustainable software development. In 2011 Annual IEEE India Conf. 1–6. https://doi.org/10.1109/INDCON.2011.6139638
@@
- Sikand et al. (2023) Samarth Sikand, Vibhu Saujanya Sharma, Vikrant Kaulgud, and Sanjay Podder. 2023. Green AI Quotient: Assessing Greenness of AI-based software and the way forward. In 38th IEEE/ACM Int. Conf. on Automated Software Engineering, ASE 2023. IEEE, 1828–1833. https://doi.org/10.1109/ASE56229.2023.00115
- Simon et al. (2023) Thibault Simon, Pierre Rust, Romain Rouvoy, and Joël Penhoat. 2023. Uncovering the Environmental Impact of Software Life Cycle. In Int. Conf. on ICT for Sustainability, ICT4S 2023, Rennes, France, June 5-9, 2023. IEEE, 176–187. https://doi.org/10.1109/ICT4S58814.2023.00026
@@
- Tee et al. (2014a) Mcxin Tee, Rusli Abdullah, and Salfarina Abdullah. 2014a. Toward Developing Green Software Development Model in Managing Knowledge of IT Practitioners for Sustaining Future Generation. In Knowledge Management Int. Conf. 2014 (KMICe).
- Tee et al. (2014b) Mcxin Tee, Rusli Abdullah, and Salfarina Abdullah. 2014b. Web-based knowledge management model for managing and sharing green knowledge of software development in community of practice. In 2014 8th. Malaysian Software Engineering Conf. (MySEC). 210–215. https://doi.org/10.1109/MySec.2014.6986016
@@
- Verdecchia et al. (2021) Roberto Verdecchia, Patricia Lago, Christof Ebert, and Carol de Vries. 2021. Green IT and Green Software. IEEE Softw. 38, 6 (2021), 7–15. https://doi.org/10.1109/MS.2021.3102254